# The Open Journal of Astrophysics

### v2

@@ 4.2 Aleatoric and Epistemic Uncertainty in IAEmu Predictions
Aleatoric uncertainty represents the intrinsic variability in the data, in this case representing variance in the correlations due to galaxy shape noise and sample variance, as studied in Van Alfen et al. (2024). The aleatoric uncertainties of $\omega$ and $\eta$ can thus be reduced through a larger simulation box size (resulting in more galaxies) and through multiple realizations of the same volume. Shape noise dominates over sample variance in the HOD model predictions on the scales considered here (Van Alfen et al., 2024), making multiple realizations important for retrieving accurate correlation functions. Epistemic uncertainties are uncertainties inherent to a model and can be large when an architecture is ill-suited for a task, or when a model is not trained on sufficient data (Hüllermeier & Waegeman, 2021). Aleatoric uncertainties are directly output from IAEmu through its design and training procedure. Epistemic uncertainties are obtained via the Monte Carlo dropout technique (Gal & Ghahramani, 2016), where dropout is used during inference across multiple forward passes. This introduces stochasticity into IAEmu’s predictions, and the resulting variance in the outputs represents the epistemic uncertainty (see Hüllermeier & Waegeman (2021) for a review on distinguishing between aleatoric and epistemic uncertainty).
@@
Figure 4, which compares the aleatoric and epistemic uncertainties from IAEmu with the true aleatoric uncertainty from halotools-IA across 10 realizations of the simulation. The figure shows that epistemic uncertainties are generally smaller than aleatoric uncertainties, as indicated by the majority of scatter points falling below the 1:1 line. This suggests that IAEmu’s architecture is sufficiently expressive for this task, and that it was not data-limited during training despite the stochasticity of these correlations. However, a median bias of 0.34 dex for $\omega$ and 0.18 dex for $\eta$ for aleatoric uncertainties when compared to the true aleatoric uncertainties is observed, suggesting that IAEmu is not perfectly calibrated for aleatoric uncertainties. This residual is particularly pronounced near the 1:1 line, wherein IAEmu’s epistemic uncertainty predictions are comparable to the aleatoric uncertainty predictions. That is, IAEmu is prone to overestimate aleatoric uncertainties when it is more uncertain of the correlation amplitudes. Nevertheless, the shape noise estimates from IAEmu can provide valuable covariance information for Monte Carlo inference (Berman et al., 2025), significantly improving posterior constraints compared to inference without covariance information.
@@
We also study the relationship between aleatoric and epistemic uncertainties in terms of the residuals in the correlation amplitudes, as shown in Figure 5. For $\omega$, we observe a trend where the largest errors in the correlation amplitudes occur in the regime where the epistemic uncertainties are 1 dex smaller than the predicted aleatoric uncertainties. This trend is more pronounced for $\eta$, where the highest errors occur when the aleatoric uncertainties are 2 dex larger than IAEmu’s epistemic uncertainties, as seen in Figure 4. This may indicate an overconfidence for IAEmu predictions of $\eta$ in this regime; however, it is also clear in comparing Figure 4 and Figure 5 that this regime is where the $\eta$ correlations are noisiest. It is thus expected that the residual on IAEmu predictions would be exaggerated due to IAEmu not overfitting to the shape noise. Nonetheless, this regime is also where IAEmu aleatoric uncertainty predictions are the most accurate.
@@
  • IAEmu is not limited by data, as evidenced by the scale of its epistemic uncertainties compared to aleatoric uncertainties for both $\omega$ and $\eta$.
@@
  • IAEmu tends to overestimate aleatoric uncertainties for both $\omega$ and $\eta$ in regimes where they are comparable to the epistemic uncertainties. This is when the model is most uncertain. IAEmu correlation amplitudes are still accurate in this regime, as shown in Figure 5.
@@
In practice, one may consider both the aleatoric and epistemic uncertainties predicted by IAEmu to assess the quality of its predictions in the absence of an underlying halotools-IA ground truth. Despite the observed bias, aleatoric uncertainty remains valuable for covariance estimation (e.g., accounting for shape noise) when performing parameter inference with IAEmu (Berman et al., 2025). Post-hoc calibration methods, such as those discussed in Grandón & Sellentin (2022), can help correct for these biases in parameter inference. Even when the primary concern is the correlation amplitudes, the relationship between IAEmu’s epistemic and aleatoric uncertainties provides valuable insight into the reliability of the predictions, as illustrated in Figures 4 and 5.
4.3 Parameter Estimation with IllustrisTNG
@@
Figure 6: The two-point correlation functions (2PCFs) for IA, fitted to observations from the TNG300 simulation, using both halotools-IA and IAEmu. These 2PCFs correspond to the posterior mean values of $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$, as shown in Figure LABEL:fig:corner. Error bars for TNG300 are obtained via jackknife resampling, while the $1\sigma$ epistemic uncertainty for IAEmu is estimated from 50 forward passes using the Monte Carlo dropout technique. The $1\sigma$ uncertainty band for halotools-IA reflects variations from random realizations of the model. Left: Position-position correlation function $\xi$ with the upper and lower curves offset by 1 dex for visual clarity, showing that IAEmu can model galaxy bias. Middle: Position-orientation correlation function $\omega$. Right: Orientation-orientation correlation function $\eta$.
@@
Our results in the previous section were obtained using a test set held out during training from the HOD simulation dataset. This demonstrates that the model generalizes well to novel data drawn from the same distribution as the training set. Previously, Van Alfen et al. (2024) showed that halotools-IA is expressive enough to model the IA signal derived from The IllustrisTNG300 suite of hydrodynamical simulations, which incorporate more complex physics, including baryonic effects. This constitutes an out-of-distribution (OOD) shift over the joint distribution of inputs and outputs from an HOD that IAEmu was trained on. In this section, we investigate whether IAEmu exhibits a similar modeling capability as halotools-IA, and can thus be robust to OOD shifts for inverse modeling. To this end, we select the best-fit occupation model parameters that reproduce the HOD of TNG300, as described in Van Alfen et al. (2024), and determine the posterior distributions on $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ that fit the signal. This is to ensure that halos with comparable masses are populated with a comparable number of galaxies as in TNG300, leaving galaxy alignment as the major factor affecting how similar correlations from the two samples are. This experiment therefore enables us to investigate potential biases between IAEmu and halotools-IA in the alignment parameter input space when modeling IA for an OOD sample. To perform parameter inference, we leverage the differentiability of IAEmu to attain efficient posterior estimates. One of the advantages of IAEmu, and neural networks in general, is its ability to act as a differentiable forward model. This property is particularly useful for inverse problems, where the goal is to perform parameter inference based on given observations. By exploiting this differentiability, one can employ a range of differentiable sampling algorithms to obtain posterior distributions for the parameters. While Van Alfen et al. (2024) used MCMC for this purpose, we instead use Hamiltonian Monte Carlo (HMC) (Duane et al., 1987) to achieve posterior estimates more efficiently by leveraging the gradient information inherent in IAEmu. Further theoretical background for HMC can be found in Appendix C.
@@
We employ Hamiltonian Monte Carlo (HMC) with the No U-Turn Sampler (NUTS, Hoffman & Gelman (2011)), using $2000$ warm-up steps and an initial learning rate of $0.005$, collecting $4000$ posterior samples for analysis. All posteriors resulted in an effective sample size greater than 1000, and all HMC experiments were executed on a single GPU and converged in roughly one minute. For comparison, the MCMC implementation in Van Alfen et al. (2024) utilized parallelization across 150 CPU cores and required up to a full day due to computational constraints. This highlights a near 2000×2000$\times$2000 × speed-up for IAEmu-HMC relative to halotools-IA-MCMC on the tested hardware. While this is somewhat lower than the acceleration achieved in forward modeling with IAEmu compared to halotools-IA, it remains a substantial improvement. The reduced gain is anticipated: HMC necessitates backpropagation through IAEmu, roughly doubling the computational load compared to a forward pass (Goodfellow et al., 2016), along with added overhead from NUTS numerical integration. Additionally, HMC’s sequential nature does not allow for parallelization. Nevertheless, its rapid convergence demonstrates its efficiency over traditional, parallelized MCMC approaches. We emphasize that a detailed convergence comparison was not performed, and that the parallelized MCMC yielded approximately $75,000$ posterior samples, in contrast to the $4000$ obtained via IAEmu. Hence, the reported speed-up metrics should be interpreted as indicative benchmarks rather than definitive measurements.
@@
The corner plots in Figure LABEL:fig:corner show the joint $(\mu_{\text{cen}},\mu_{\text{sat}})$ posteriors for three separate stellar mass thresholds $M_{*}$ for both IAEmu and halotools-IA. Sample 1 corresponds to $\log(M_{*})>10.5$, Sample 2 to $\log(M_{*})>10.0$, and Sample 3 to $\log(M_{*})>9.5$. The HOD parameter fits corresponding to these mass cutoffs can be found in (Van Alfen et al., 2024). We confirm two trends also observed in Van Alfen et al. (2024): central alignment strength is larger than satellite alignment strength, and the alignment strength monotonically increases with the stellar mass threshold. We find a greater than $0.4\sigma$ agreement between MCMC with halotools-IA and HMC with IAEmu for all samples. The strongest discrepancy is in the posterior variance for Sample 3, which is the noisiest set of Illustris data and also has the largest IAEmu epistemic uncertainty, as seen in Figure 6. This reflects the discussion in Section 4.2, wherein it was seen that the epistemic uncertainty of IAEmu is correlated with the true and predicted aleatoric uncertainty. Exact values for halotools-IA and IAEmu posterior summary statistics are shown in Table 1.
@@
halotools-IA $0.54^{+0.04}_{-0.04}$ $0.05^{+0.03}_{-0.03}$
The correlation function predictions from IAEmu with posterior means for $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ are shown in Figure 6, in which we see that there is generally good agreement with IAEmu predictions compared to those from halotools-IA for all correlations. The correlation function, $\xi$, is also shown to illustrate the agreement between halotools-IA and IAEmu for galaxy clustering statistics; however, it does not depend on $\mu_{\text{cen}}$ or $\mu_{\text{sat}}$.
@@ 5 Summary & Discussion
In this work, we have developed a neural network-based surrogate model, IAEmu, designed to predict galaxy intrinsic alignment correlations derived from halo occupation distribution modeling. IAEmu eliminates the need to generate full galaxy catalogs and compute correlation functions using traditional HOD pipelines, which are computationally expensive. On a single GPU, IAEmu achieves a ×$\bm{\times 10^{4}}$ speed-up in wall-clock time compared to halotools-IA run on a moderately parallelized CPU setup representative of typical resources (e.g., $\sim$150 cores). When comparing single GPU to single CPU performance, this corresponds to an approximate $10^{6}\times$ speed-up. This substantial acceleration enables efficient forward modeling and significantly expedites inverse modeling tasks. The differentiable nature of IAEmu facilitates the use of gradient-based inference methods such as Hamiltonian Monte Carlo (HMC), which are otherwise infeasible with halotools-IA. Although the speed-up in individual evaluations is dramatic, the end-to-end improvement in sampling-based inference compared to parallelized MCMC is somewhat lower, due to the additional computational overhead from gradient evaluations and the inherently sequential nature of HMC arising from numerical trajectory integration. Nevertheless, HMC achieves significantly faster convergence than parallelized MCMC, making it a far more efficient option overall for inverse modeling despite the reduced relative speed-up.
@@
Simultaneously training on three distinct correlations, each with varying scales and signal-to-noise, presented additional engineering challenges from a training perspective. In particular, the low signal-to-noise of the $\eta$ correlation made accurate modeling of its uncertainties especially difficult. To address this, we implemented the $\beta$-NLL loss within a multitask learning framework. This approach ensured that correlations with larger amplitudes did not disproportionately dominate the loss landscape, while facilitating effectively individualized training for each decoder branch, allowing us to prioritize more accurate aleatoric uncertainty estimates for the noisier $\omega$ and $\eta$ correlations while focusing on correlation amplitude predictions for the higher-signal $\xi$ correlation.
@@
Finally, we found that IAEmu has comparable performance to halotools-IA when used to fit the alignment parameters $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ to IA correlation measurements from the TNG300 hydrodynamic simulation, in a manner similar to the robustness test originally performed in (Van Alfen et al., 2024). This demonstrates IAEmu’s robustness to OOD shifts for inverse problems. Specifically, we observe a better than $0.4\sigma$ agreement in the $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ posteriors across three separate mass regimes between IAEmu, fit using HMC, and halotools-IA, fit with Markov Chain Monte Carlo (MCMC). A significant advantage was the improvement in computational efficiency; while halotools-IA with MCMC required approximately one day on a cluster CPU, IAEmu completed the inverse problem in less than a minute on a single GPU. This constitutes a nearly 2000x speed up over MCMC with halotools-IA, demonstrating that the efficiency benefits of neural network surrogate models extend beyond forward modeling.
@@
Future Work. In future work, we will improve upon the IAEmu pipeline by exploring different architectural and data choices. IAEmu, as presented here, operates in a traditional supervised learning regime, where the model learns a direct, deterministic mapping between the HOD parameters and correlations. Although we introduce stochasticity for uncertainty quantification via MC dropout and MVE, a more natural probabilistic approach could be achieved through conditional diffusion generative modeling, where the model learns a probabilistic mapping via a denoising process on the data, or through flow-based architectures as in Pandey et al. (2024). These models can also be made symmetry-aware, enhancing their effectiveness in physical settings like this (see Jagvaral et al., 2023a, for an example of SO(3)-equivariant diffusion applied to IA). Diffusion models typically leverage architectures like U-Nets (Ronneberger et al., 2015), which are well-suited for denoising tasks (Schanz et al., 2023) and follow an encoder-decoder architecture similar to what was used here. The denoising training paradigm when applied to cosmology has thus far exhibited promising results in enhancing the resolution of existing simulations (Schanz et al., 2023) as well as functioning as surrogate models (Mudur et al., 2024). Other avenues for multi-task learning, such as treating the loss coefficients as trainable parameters as in (Kendall et al., 2018), can allow more effective optimization without normalization and can potentially alleviate the biased performance in $\xi$ observed here.
Furthermore, these future avenues would open opportunities in studying the variation in cosmological parameters, a dimension not explored in this work. Here, we have only used a single underlying dark matter catalog from an N-body simulation, which implicitly reflects specific cosmological parameters. It would be exciting to explore IA modeling for varying dark matter catalogs and thus different cosmologies. In a pure HOD setting, these extensions can be expensive.
@@
Simulation-based modeling has led to tremendous progress in understanding galaxy intrinsic alignments, particularly when compared to earlier analytic and semi-analytic methods. However, this progress has traditionally come with incurred computational expense. In this work, we have shown a compelling case in which accuracy and efficiency can be achieved with NN-based emulators for galaxy intrinsic alignments from HOD simulations. This is a significant step towards accelerating model validation strategies in preparation for data from the Rubin Observatory, Roman Space Telescope, and other Stage IV surveys.

### v4

@@ 4.2 Aleatoric and Epistemic Uncertainty in IAEmu Predictions
Aleatoric uncertainty represents the intrinsic variability in the data, in this case representing variance in the correlations due to galaxy shape noise and sample variance, as studied in vanalfen_2023. The aleatoric uncertainties of $\omega$ and $\eta$ can thus be reduced through a larger simulation box size (resulting in more galaxies) and through multiple realizations of the same volume. Shape noise dominates over sample variance in the HOD model predictions (vanalfen_2023), making multiple realizations important for retrieving accurate correlation functions. Epistemic uncertainties are uncertainties inherent to a model and can be large when an architecture is ill-suited for a task, or when a model is not trained on sufficient data (Hullermeier_2021). Aleatoric uncertainties are directly output from IAEmu through its design and training procedure. Epistemic uncertainties are obtained via the Monte Carlo dropout technique (gal2016dropout), where dropout is used during inference across multiple forward passes. This introduces stochasticity into IAEmu’s predictions, and the resulting variance in the outputs represents the epistemic uncertainty (see Hullermeier_2021 for a review on distinguishing between aleatoric and epistemic uncertainty).
@@
Figure 4 compares the aleatoric and epistemic uncertainties from IAEmu with the true aleatoric uncertainty from halotools-IA across 10 realizations of the simulation. The figure shows that epistemic uncertainties are generally smaller than aleatoric uncertainties, as indicated by the majority of scatter points falling below the 1:1 line. This suggests that IAEmu’s architecture is sufficiently expressive for this task, and that it was not data-limited during training despite the stochasticity of these correlations. However, a median bias of 0.34 dex for $\omega$ and 0.18 dex for $\eta$ for aleatoric uncertainties when compared to the true aleatoric uncertainties is observed, suggesting that IAEmu is not perfectly calibrated for aleatoric uncertainties. This residual is particularly pronounced near the 1:1 line, wherein IAEmu’s epistemic uncertainty predictions are comparable to the aleatoric uncertainty predictions. That is, IAEmu tends to overestimate aleatoric uncertainties in regions where the correlation amplitudes are less certain. Nevertheless, the shape noise estimates from IAEmu can provide valuable covariance information for Monte Carlo inference (berman2025softclustering), significantly improving posterior constraints compared to inference without covariance information.
@@
We also study the relationship between aleatoric and epistemic uncertainties in terms of the residuals in the correlation amplitudes, as shown in Figure 5. For $\omega$, we observe a trend where the largest errors in the correlation amplitudes occur in the regime where the epistemic uncertainties are 1 dex smaller than the predicted aleatoric uncertainties. This trend is more pronounced for $\eta$, where the highest errors occur when the aleatoric uncertainties are 2 dex larger than IAEmu’s epistemic uncertainties, as seen in Figure 4. This may indicate an overconfidence for IAEmu predictions of $\eta$ in this regime; however, it is also clear in comparing Figure 4 and Figure 5 that this regime is where the $\eta$ correlations are the noisiest. In other words, this occurs when the galaxy shape noise is most prominent. It is thus expected that the residual on IAEmu predictions would be exaggerated due to IAEmu not overfitting to the shape noise. Nonetheless, this regime is also where IAEmu aleatoric uncertainty predictions are the most accurate.
@@
  • IAEmu is not significantly limited by data, as evidenced by the scale of its epistemic uncertainties compared to aleatoric uncertainties for $\omega$ and $\eta$.
@@
  • IAEmu tends to overestimate aleatoric uncertainties for both $\omega$ and $\eta$ in regimes where they are comparable to the epistemic uncertainties. This occurs when the model is most uncertain. IAEmu correlation amplitudes are still accurate in this regime, as shown in Figure 5.
@@
In practice, one may consider both the aleatoric and epistemic uncertainties predicted by IAEmu to assess the quality of its predictions in the absence of an underlying halotools-IA ground truth. Despite the observed bias, aleatoric uncertainty remains valuable for covariance estimation (e.g., accounting for shape noise) when performing parameter inference with IAEmu (berman2025softclustering). Post-hoc calibration methods, such as those discussed in Grand_n_2022, can help correct for these biases in parameter inference. These methods calibrate uncertainty estimates after training, ranging from non-parametric approaches like histogram binning and isotonic regression, to simple parametric schemes such as Platt scaling (guo2017calibrationmodernneuralnetworks) and its extensions (e.g. temperature scaling and Dirichlet calibration). Even when the primary concern is the correlation amplitudes, the relationship between IAEmu’s epistemic and aleatoric uncertainties provides valuable insight into the reliability of the predictions, as illustrated in Figures 4 and 5.
4.3 Parameter Estimation with tng300
@@
Figure 6: The two-point correlation functions (2PCFs) for IA, fitted to observations from the tng300 simulation, using both halotools-IA and IAEmu. The correlations are measured across three mass threshold samples, as denoted in the left panel legend. Purple corresponds to most massive sample, pink for intermediate, and red for least massive. True correlations are shown as scatter points and HOD and IAEmu fits shown as lines. These 2PCFs correspond to the posterior mean values of $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$, as shown in Figure 9(a). Error bars for tng300 are obtained via jackknife resampling, while the $1\sigma$ epistemic uncertainty for IAEmu is estimated from 50 forward passes using the Monte Carlo dropout technique. The $1\sigma$ uncertainty band for halotools-IA reflects variations from random realizations of the model. Left: Position-position correlation function $\xi$ with the upper and lower curves offset by 1 dex for visual clarity, showing that IAEmu can model galaxy bias. Middle: Position-orientation correlation function $\omega$. Right: Orientation-orientation correlation function $\eta$.
@@
Our results in the previous section were obtained using a test set held out during training from the HOD simulation dataset. This demonstrates that the model generalizes well to novel data drawn from the same distribution as the training set. Previously, vanalfen_2023 showed that halotools-IA is expressive enough to model the IA signal derived from The tng300 suite of hydrodynamical simulations, which incorporate more complex physics, including baryonic effects. This constitutes an out-of-distribution (OOD) shift over the joint distribution of inputs and outputs from an HOD that IAEmu was trained on. In this section, we investigate whether IAEmu exhibits a similar modeling capability as halotools-IA, and can thus be robust to OOD shifts for inverse modeling. To this end, we select the best-fit occupation model parameters that reproduce the HOD of tng300, as described in vanalfen_2023, and determine the posterior distributions on $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ that fit the signal. This ensures that halos with comparable masses are populated with a comparable number of galaxies as in tng300, leaving galaxy alignment as the major factor affecting how similar correlations from the two samples are. This experiment therefore enables us to investigate potential biases between IAEmu and halotools-IA in the alignment parameter input space when modeling IA for an OOD sample.
To perform parameter inference, we leverage the differentiability of IAEmu to attain efficient posterior estimates. One of the advantages of IAEmu, and neural networks in general, is its ability to act as a differentiable forward model. This property is particularly useful for inverse problems, where the goal is to perform parameter inference based on given observations. By exploiting this differentiability, one can employ a range of differentiable sampling algorithms to obtain posterior distributions for the parameters. While vanalfen_2023 used MCMC for this purpose, we instead use Hamiltonian Monte Carlo (HMC) (DUANE1987216) to achieve posterior estimates more efficiently by leveraging the gradient information inherent in IAEmu. Further theoretical background for HMC can be found in Appendix C.
@@
We employ Hamiltonian Monte Carlo (HMC) with the No U-Turn Sampler (NUTS, hoffman2011nouturnsampleradaptivelysetting), using $2000$ warm-up steps and an initial learning rate of $0.005$, collecting $4000$ posterior samples for analysis. All posteriors resulted in an effective sample size greater than 1000, and all HMC experiments were executed on a single GPU and converged in roughly one minute. For comparison, the MCMC implementation in vanalfen_2023 utilized parallelization across 150 CPU cores and required up to a full day due to computational constraints. This highlights a near $2000\times$ speed-up for IAEmu-HMC relative to halotools-IA-MCMC on the tested hardware. While this is somewhat lower than the acceleration achieved in forward modeling with IAEmu compared to halotools-IA, it remains a substantial improvement. The reduced gain is anticipated: HMC necessitates backpropagation through IAEmu, roughly doubling the computational load compared to a forward pass (Goodfellow-et-al-2016), along with added overhead from NUTS numerical integration. Additionally, the sequential nature of HMC does not allow for parallelization. Nevertheless, its rapid convergence demonstrates its efficiency over traditional, parallelized MCMC approaches. We emphasize that a detailed convergence comparison was not performed, and that the parallelized MCMC yielded approximately $75,000$ posterior samples, in contrast to the $4000$ obtained via IAEmu. Hence, the reported speed-up metrics should be interpreted as indicative benchmarks rather than definitive measurements.
(a) Optimal parameter values for central alignment strength ($\mu_{\text{cen}}$) and satellite alignment strength ($\mu_{\text{sat}}$) fit to $\omega$ observations from tng300 with three distinct mass cutoffs for halos included in the underlying HOD model. Posterior contours for halotools-IA and IAEmu are shown with 4000 posterior samples each, with halotools-IA contours in orange and IAEmu contours in blue. Posteriors for halotools-IA were obtained via MCMC using 75 walkers running in parallel for 23 hours on CPU, resulting in up to 1300 steps per walker, or as few as about 450 steps per walker for slower runs. Posteriors for IAEmu were retrieved using NUTS, a variant of the HMC algorithm, with 2000 warm-up steps around a minute on a single GPU. IAEmu posteriors exhibit a better than $0.4\sigma$ overlap with posteriors from halotools-IA, indicating that IAEmu can generalize to OOD shifts for inverse modeling. Exact posterior summaries for comparison can be found in Table 2. Left: Sample 1 IAEmu posteriors with optimal values $\mu_{\text{cen}}=0.81$ and $\mu_{\text{sat}}=0.35$. Middle: Sample 2 IAEmu posteriors with optimal values $\mu_{\text{cen}}=0.70$ and $\mu_{\text{sat}}=0.14$. Right: Sample 3 IAEmu posteriors with optimal values $\mu_{\text{cen}}=0.52$ and $\mu_{\text{sat}}=0.01$.
The correlation function predictions from IAEmu with posterior means for $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ are shown in Figure 6, in which we see that there is generally good agreement with IAEmu predictions compared to those from halotools-IA for all correlations. The correlation function, $\xi$, is also shown to illustrate the agreement between halotools-IA and IAEmu for galaxy clustering statistics; however, it does not depend on $\mu_{\text{cen}}$ or $\mu_{\text{sat}}$. We also see that the quality of fit for both halotools-IA and IAEmu decreases with decreasing stellar mass, which was also observed in vanalfen_2023. This provides some indication that a constant alignment strength parameterization is not sufficient to extend this HOD-based model to hydrodynamic simulations in the low-mass regime, which could potentially be addressed with a mass-dependent alignment parameterization.
@@
The corner plots in Figure 9(a) show the joint $(\mu_{\text{cen}},\mu_{\text{sat}})$ posteriors for three separate stellar mass thresholds $M_{*}$ for both IAEmu and halotools-IA. Sample 1 corresponds to $\log(M_{*})>10.5$, Sample 2 to $\log(M_{*})>10.0$, and Sample 3 to $\log(M_{*})>9.5$. The HOD parameter fits corresponding to these mass cutoffs can be found in (vanalfen_2023). We confirm two trends also observed in vanalfen_2023: central alignment strength is larger than satellite alignment strength, and the alignment strength monotonically increases with the stellar mass threshold. We find a greater than $0.4\sigma$ agreement between MCMC with halotools-IA and HMC with IAEmu for all samples. The strongest discrepancy is in the posterior variance for Sample 3, which is the noisiest set of tng300 data and also has the largest IAEmu epistemic uncertainty, as seen in Figure 6. This reflects the discussion in Section 4.2, wherein it was seen that the epistemic uncertainty of IAEmu is correlated with the true and predicted aleatoric uncertainty. Exact values for halotools-IA and IAEmu posterior summary statistics are shown in Table 2.
@@
halotools-IA $54^{+0.04}_{-0.04} 0.0$$05^{+0.03}_{-0.03}$
@@ 5 Summary & Discussion
In this work, we have developed a neural network-based surrogate model, IAEmu, designed to predict galaxy intrinsic alignment correlations derived from halo occupation distribution modeling. IAEmu eliminates the need to generate full galaxy catalogs and computes correlation functions using traditional HOD pipelines, which are computationally expensive. On a single GPU, IAEmu achieves a $\bm{\times 10^{4}}$ speed-up in wall-clock time compared to halotools-IA run on a moderately parallelized CPU setup representative of typical resources (e.g., $\sim$150 cores). When comparing single GPU to single CPU performance, this corresponds to an approximate $10^{6}\times$ speed-up. This substantial acceleration enables efficient forward modeling and significantly expedites inverse modeling tasks. The differentiable nature of IAEmu facilitates the use of gradient-based inference methods such as Hamiltonian Monte Carlo (HMC), which are otherwise infeasible with halotools-IA. Although the speed-up in individual evaluations is dramatic, the end-to-end improvement in sampling-based inference compared to parallelized MCMC is somewhat lower, due to the additional computational overhead from gradient evaluations and the inherently sequential nature of HMC arising from numerical trajectory integration. Nevertheless, HMC achieves significantly faster convergence than parallelized MCMC, making it a far more efficient option overall for inverse modeling despite the reduced relative speed-up.
@@
Simultaneously training on three distinct correlations, each with varying numerical scales and signal-to-noise presented additional engineering challenges from a training perspective. In particular, the low signal-to-noise of the $\eta$ correlation made accurate modeling of its uncertainties especially difficult. To address this, we implemented the $\beta$-NLL loss within a multitask learning framework. This approach ensured that correlations with larger amplitudes did not disproportionately dominate the loss landscape, while facilitating effectively individualized training for each decoder branch, allowing us to prioritize more accurate aleatoric uncertainty estimates for the noisier $\omega$ and $\eta$ correlations while focusing on correlation amplitude predictions for the higher-signal $\xi$ correlation.
@@
Finally, we found that IAEmu has comparable performance to halotools-IA when used to fit the alignment parameters $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ to IA correlation measurements from the tng300 hydrodynamic simulation, in a manner similar to the robustness test originally performed in (vanalfen_2023). This demonstrates IAEmu’s robustness to OOD shifts for inverse problems. Specifically, we observe a better than $0.4\sigma$ agreement in the $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ posteriors across three separate mass regimes between IAEmu, fit using HMC, and halotools-IA, fit with Markov Chain Monte Carlo (MCMC). A significant advantage was the improvement in computational efficiency; while halotools-IA with MCMC required approximately one day on a cluster CPU, IAEmu completed the inverse problem in less than a minute on a single GPU. This constitutes a nearly $2000\times$ speed up over MCMC with halotools-IA, demonstrating that the efficiency benefits of neural network surrogate models extend beyond forward modeling.
Limitations. IAEmu is an emulator designed to predict 2PCFs based on HOD modeling. Consequently, one inherent limitation of IAEmu is its reliance on phenomenological HOD parameterizations, and IAEmu thus inherits any limitations currently present in halotools-IA. Moreover, since IAEmu was trained solely on HODs conducted on the Bolshoi-Planck $N$-body simulation, it does not currently factor in cosmology dependence, which is an avenue of future work.
Future Work. There remain several functional improvements required for IAEmu to be fully deployable in cosmological analyses. At present, IAEmu predicts only 2PCFs and is not configured to model corresponding one-point statistics such as galaxy number counts. Furthermore, the 2PCFs computed in halotools-IA only include galaxy orientations and not their full shapes, which are essential for a complete IA analysis. In addition, emerging simulation suites are beginning to forego explicit host–subhalo distinctions in favor of halo-core models (vitório2025exploringcoregalaxyconnection). As a result, halotools-IA would need to be updated to incorporate halo-core models, and IAEmu would accordingly need to be retrained. Incorporating these improvements is ongoing work in both halotools-IA and future iterations of IAEmu. Lastly, IAEmu also has the capacity to perform joint inference over both HOD and IA parameters; however, this was not explored in Section 4.3. The central purpose of this experiment was reproducing the results of vanalfen_2023, which focused on varying the IA parameters. Performing a joint inference over both HOD and IA parameters will be explored in future iterations of IAEmu.
@@
In future versions of the emulator, we will improve upon the IAEmu pipeline by exploring different architectural, data, and modeling choices. IAEmu, as presented here, operates in a traditional supervised learning regime, where the model learns a direct, deterministic mapping between the HOD parameters and correlations. Although we introduce stochasticity for uncertainty quantification via MC dropout and MVE, a more natural probabilistic approach could be achieved through conditional diffusion generative modeling, where the model learns a probabilistic mapping via a denoising process on the data, or through flow-based architectures as in pandey2024charmcreatinghalosautoregressive. These models can also have their internal representations restricted by known symmetries in the data, enhancing their effectiveness in physical settings like this (see jagvaral2023diffusion, for an example of $SO(3)$-equivariant diffusion applied to IA), which naturally lends such techniques to field-level modeling of the full galaxy catalog. Field level emulation can also expand IAEmu summary statistics outside the 2PCFs modeled here. The denoising training paradigm when applied to cosmology has thus far exhibited promising results in enhancing the resolution of existing simulations (schanz2023stochasticsuperresolutioncosmologicalsimulations) as well as functioning as surrogate models (mudur2024diffusionhmcparameterinferencediffusion).
A field-level emulator for halotools-IA could incorporate elements of NN–based modeling together with differentiable components, such as differentiable HOD models (horowitz2022differentiablestochastichalooccupation). Both NNs and differentiable simulations share a differentiable structure, allowing their components to be integrated within the same modeling framework. While non-NN-based differentiable methods provide useful tools for inverse modeling, they can be computationally demanding and are often complemented by faster NN-based surrogates.
@@
Simulation-based modeling has opened a new set of opportunities to better understand galaxy intrinsic alignments, complementing earlier analytic and semi-analytic efforts. However, these new techniques have incurred additional computational expense. In this work, we have shown a compelling case in which accuracy and efficiency can be achieved with NN-based emulators for galaxy intrinsic alignments from HOD simulations. This provides a significant step towards accelerating model validation strategies in preparation for data from the Rubin Observatory, Roman Space Telescope, and other Stage IV surveys.